### v1

# Mechanistic Interpretability of LoRA-Adapted Language Models for Nuclear Reactor Safety Applications

Abstract
The integration of Large Language Models (LLMs) into safety-critical domains, such as nuclear engineering, necessitates a deep understanding of their internal reasoning processes. This paper presents a novel methodology for interpreting how an LLM encodes and utilizes domain-specific knowledge, using a Boiling Water Reactor system as a case study. We adapted a general-purpose LLM (Gemma-3-1b-it) to the nuclear domain using a parameter-efficient fine-tuning technique known as Low-Rank Adaptation. By comparing the neuron activation patterns of the base model to those of the fine-tuned model, we identified a sparse set of neurons whose behavior was significantly altered during the adaptation process. To probe the causal role of these specialized neurons, we employed a neuron silencing technique. Our results demonstrate that while silencing most of these specialized neurons individually did not produce a statistically significant effect, deactivating the entire group collectively led to a statistically significant degradation in task performance. Qualitative analysis further revealed that silencing these neurons impaired the model’s ability to generate detailed, contextually accurate technical information. This paper provides a concrete methodology for enhancing the transparency of an opaque black-box model, allowing domain expertise to be traced to verifiable neural circuits. This offers a pathway towards achieving nuclear-grade artificial intelligence (AI) assurance, addressing the verification and validation challenges mandated by nuclear regulatory frameworks (e.g., 10 CFR 50 Appendix B), which have limited AI deployment in safety-critical nuclear operations.
[…]
Yoon Pyo Leea
[…]
lukeyounpyo@hanyang.ac.kr
[…]
aDepartment of Nuclear Engineering, Hanyang University, Seoul, Republic of Korea
[…]
1 INTRODUCTION
The nuclear industry operates under a stringent regulatory and safety culture in which every component and decision-making tool must be transparent, auditable, and verifiable [1]. Foundational principles, such as defense-in-depth, which requires multiple independent layers of protection, and As Low As Reasonably Achievable, which mandates minimizing all risks [2], demand a high level of assurance for any new technology introduced into plant operations. Recent advances in Large Language Models (LLMs), built upon the Transformer architecture [3], offer the potential for knowledge management and operational support [4, 5]. However, their deployment in safety-critical environments faces a fundamental barrier: the black-box nature of their reasoning processes, which is in conflict with the industry’s core safety principles.
The deployment of software in safety-critical systems within a nuclear power plant must adhere to stringent quality assurance (QA) requirements, which is fundamentally different from its use in non-critical applications. In the United States (U.S.), regulations such as 10 CFR 50, Appendix B, mandate a documented design basis, comprehensive testing, and full traceability of a system’s logic and behavior [6]. This regulatory framework presents a significant challenge for advanced artificial intelligence (AI) and LLMs. The inherent black-box nature of these models makes it difficult, if not impossible, to verify the internal reasoning path an LLM took to arrive at a recommendation, conflicting with QA standards. This has largely limited AI deployment in nuclear facilities for non-safety roles. Therefore, a new paradigm focusing on internal transparency is required to bridge the gap between the capabilities of modern AI and the safety standards of the nuclear industry.
This paper addresses this gap by proposing a methodology based on mechanistic interpretability to enhance the trustworthiness of LLMs in nuclear applications. The goal is to increase the transparency of these opaque models, making their internal reasoning processes examinable and validatable, a prerequisite for qualification in systems important to safety, as outlined in key industrial standards such as the Institute of Electrical and Electronics Engineers (IEEE) Std 7-4.3.2 [7], and their associated regulatory guidelines, such as the U.S. National Regulatory Commission (NRC) Regulatory Guide 1.152 [8]. Recent initiatives like the NRC’s Technology-Inclusive Regulatory Framework [9] recognize the potential benefits of advanced digital systems while emphasizing the need for robust safety assurance methodologies. We demonstrate this approach by adapting a general-purpose LLM [10] to the specific domain of Boiling Water Reactors (BWRs) using documentation from authoritative bodies like the International Atomic Energy Agency (IAEA) and Organisation for Economic Co-operation and Development (OECD)/Nuclear Energy Agency (NEA) [11, 12].
To perform this domain adaptation efficiently without the prohibitive cost of full-model retraining, we leverage Parameter-Efficient Fine-Tuning (PEFT) methods. Although various techniques exist, such as Prefix-Tuning [13] and Prompt-Tuning [14], our study employed Low-Rank Adaptation (LoRA) [15], a prominent technique known for its effectiveness in specialized tasks [16]. However, verification remains challenging, even after successful adaptation. The trust gap in AI, where models may exhibit unreliable behavior or hidden biases [17], must be closed before such tools can be used in nuclear operations.
The integration of advanced digital systems into nuclear operations is an ongoing effort in the industry. Previous studies have addressed the modernization of instrumentation and control systems [18], human factors considerations for AI applications in nuclear plants [19, 20], and software quality assurance for nuclear computational tools [21]. These foundational studies established the importance of verification and validation (V&V) processes for safety-critical nuclear software systems, underscoring the need for interpretable AI methods in nuclear applications.
Therefore, our research moves beyond simple fine-tuning to investigate the internal mechanisms of the adapted model. We were guided by the following objectives:
  1. 1.
    Identify specific neurons whose activation patterns are systematically altered by LoRA fine-tuning on a corpus of BWR technical data.
  2. 2.
    Use causal interventions to test the functional importance of these identified neurons, with performance quantitatively measured using the established Bilingual Evaluation Understudy (BLEU) [22] metrics.
  3. 3.
    Propose a verifiable, step-by-step methodology for tracing abstract knowledge to the physical substrate within a neural network, thereby enhancing its interpretability.
By demonstrating that domain expertise can be traced to a small, verifiable set of neural components, this study provides a pathway towards achieving nuclear-grade AI assurance. This mechanistic insight is a foundational step towards building the auditable and robust AI systems required for the future of nuclear science and engineering. The remainder of this paper is organized as follows. Section 2 reviews the key AI concepts relevant to this study. Section 3 describes our experimental approach, including dataset construction, model adaptation, and our causal analysis techniques. Section 4 presents the findings of the neuron silencing experiments. Section 5 interprets these results and discusses their implications for AI in terms of nuclear safety. Finally, Section 6 summarizes our work and suggests directions for future research.
2 BACKGROUND: AI, INTERPRETABILITY, AND NUCLEAR SAFETY
This section provides an overview of the key concepts underlying this study. First, it presents the central challenge of using AI in safety-critical domains, and then introduces the foundational AI techniques and interpretability methods employed to address this challenge.
2.1 Domain Adaptation of Large Language Models
An LLM is an AI model, built upon the Transformer architecture, that is pre-trained on vast text datasets to acquire general language and world knowledge [3, 23]. To adapt such a generalist model to a specialized field, a process called fine-tuning is used, where the model is further trained on a smaller, domain-specific dataset—in this case, a corpus of BWR technical documents.
However, fine-tuning all of the billions of parameters of the model is not only computationally prohibitive but also risks inducing ”catastrophic forgetting,” where its valuable pre-trained knowledge is overwritten by new, domain-specific data. We employ LoRA, a prominent PEFT technique that freezes the original model weights and injects small trainable adaptation modules (low-rank matrices) into the Transformer layers [15]. This allows for efficient domain adaptation while preserving the model’s valuable pre-trained knowledge.
2.2 Mechanistic Interpretability for Nuclear-Grade Assurance
To meet nuclear-grade assurance requirements, we must move beyond ”post-hoc” Explainable AI (XAI) techniques like Local Interpretable Model-agnostic Explanations (LIME) or SHapley Additive Explanations (SHAP), which explain predictions by correlating inputs and outputs without revealing the model’s internal workings [24, 25]. In contrast, ”mechanistic interpretability,” the approach adopted in this study, aims to directly analyze and causally verify the internal computational structures— neural circuits—that generate a model’s behavior.
In high-reliability fields such as nuclear engineering, verifying how a conclusion was reached is as important as the conclusion itself. Neuronal activation was used as a direct signal for this purpose. A neuron is a fundamental computational unit in the model, and its activation is a numerical value representing its output in response to an input. By tracking which neurons’ activations are systematically altered during domain adaptation, we can begin to trace abstract knowledge to a physical, verifiable substrate within the model, a prerequisite for satisfying the stringent V&V demands of nuclear regulators.
2.3 Language Generation Evaluation Metric
In our evaluation, we focused our quantitative analysis on the BLEU score. Given the safety-critical nature of the nuclear domain, the precision, fluency, and linguistic fidelity of the model’s output are considered paramount. An answer that is factually complete but poorly or ambiguously can lead to operational errors. Because BLEU directly measures these characteristics by evaluating n-gram precision, it was selected as the most appropriate metric for validating the model’s performance in this specific context.
3 METHODOLOGY
Our investigation into the internal mechanisms of domain adaptation followed a three-stage experimental framework: (1) construction of a domain-specific dataset, (2) PEFT of the base model, and (3) neuron-level analysis through activation tracking and causal intervention.
3.1 Dataset Construction
We constructed a bespoke Question-Answering dataset to train and evaluate the model in a realistic nuclear engineering context. The source material for this dataset was a collection of authoritative technical documents covering BWR systems, safety procedures, and simulator specifications from various international bodies. Specifically, we utilized documentation from the OECD NEA on severe accident mitigation [12], an IAEA report on BWR simulators [11], and system descriptions from the U.S. NRC Technical Training Center [26].
The Question-Answering pairs were generated using a custom Python script that leveraged OpenAI’s gpt-4o-2024-08-06 application programming interface (API) [27]. This process involved segmenting the source texts into manageable chunks and prompting the model to create relevant questions and their corresponding answers based exclusively on the text provided. The resulting dataset was divided using stratified sampling to ensure a balanced representation across the source documents in both the training and evaluation sets.
3.2 Model Fine-Tuning
The base model for our study was Gemma-3-1b-it, a state-of-the-art, general-purpose LLM [10]. For domain adaptation, we fine-tuned this model on the training split of our custom Question-Answering dataset using the LoRA [15] technique, as previously described. The LoRA configuration was applied to a comprehensive set of target modules within the model’s attention and feed-forward layers: q proj, k proj, v proj, o proj, gate proj, up proj, and down proj. The complete set of hyperparameters used in this process is listed in Table 1. This process resulted in the specialized model, referred to as ”LoRA” in our experiments.
Table 1: Fine-tuning Hyperparameters
Parameter Value
Model & Quantization
Base Model Gemma-3-1b-it
Compute Data Type bfloat16
Precision Type FP16 (for training)
LoRA Configuration
Rank (r) 8
Alpha (alpha) 16
Dropout 0.05
Target Modules q proj, k proj, v proj, o proj,
gate proj, up proj, down proj
Training Arguments
Optimizer AdamW (PyTorch)
Learning Rate 2e-5
Batch Size (per device) 1
Gradient Accum. Steps 8
Effective Batch Size 8
Epochs 2
Warmup Steps 50
3.3 Neuron-Level Analysis
The core of our analysis involved identifying and probing the neurons responsible for the model’s newly acquired domain expertise. To facilitate the extraction of internal activation and the implementation of model interventions, we utilized PyTorch’s hook mechanism for fine-grained control over neuron activation. This approach aligns with the principles of mechanistic interpretability, which allows us to probe specific neural circuits.
3.3.1 Identifying Key Neurons
We first measured the average activation of each neuron in the final Multilayer Perceptron (MLP) layer for both the base Gemma model and our fine-tuned LoRA model across the entire evaluation dataset. The change in average activation ($\Delta_{\text{act}}=\text{AvgAct}_{\text{LoRA}}-\text{AvgAct}_{\text{Base}}$) was calculated for every neuron. Neurons exhibiting the largest absolute change were designated as key neurons for further analysis. Our study focused on the top six neurons with the largest absolute change in activation, a group that comprised five amplified (positive change) and one suppressed (negative change) neuron.
3.3.2 Causal Intervention via Neuron Silencing
To test the functional role of these key neurons, we employed a causal intervention technique known as neuron silencing. Silencing was implemented by attaching a forward hook to the final MLP layer of the model using PyTorch’s register forward hook function. This hook intercepts the layer’s output tensor during the forward pass of the model and sets the activation values at the specified neuron indices to zero before they are passed to subsequent layers. This method allows for targeted, real-time intervention without altering the underlying weights of the model. For each key neuron identified, we created a modified version of the LoRA model by implementing this silencing hook. Additionally, we created a model where all identified key neurons were silenced simultaneously (”LoRA-Silenced-Key6”). The effects of these interventions were evaluated using the BLEU metric.
3.3.3 Statistical Analysis
To determine whether the observed differences in the performance metric (BLEU) between the model pairs were statistically significant, we employed the Wilcoxon signed-rank test [28]. This non-parametric test is suitable for comparing paired samples, such as the lists of scores generated by two different models on the same set of questions. A p-value of less than 0.05 was considered the threshold for statistical significance.
4 RESULTS
This section presents the experimental results. It begins with the identification of neurons affected by fine-tuning, followed by the quantitative outcomes of the neuron-silencing interventions.
4.1 LoRA Fine-Tuning Induces Neuron Specialization
Our initial analysis focused on identifying neurons affected by the LoRA fine-tuning process. Figure 1 illustrates the average change in activation of the 12 neurons that exhibited the largest absolute shifts between the base and the LoRA-tuned models. A distinct pattern emerged: one set of neurons was consistently amplified (e.g., Neurons #1086, #560, #1066, #176, and #890), showing a large positive change in activation, while another set was suppressed (e.g., Neuron #941), showing a negative change. The full activation distributions for these neurons are presented as boxplots in Figure 2, which show a consistent upward and downward shifts in the median activation values for the amplified and suppressed neurons, respectively.
Figure 1: Average activation change for the top 12 most affected neurons after LoRA fine-tuning. Positive bars indicate amplification (increased activation), while negative bars indicate suppression (decreased activation).
Figure 2: Boxplot comparison of activation values for the top 12 neurons, illustrating the distributional shift between the base and LoRA models.
4.2 Quantitative Effects of Neuron Silencing
Neuron silencing experiments were performed to measure the causal impact of the identified key neurons. Figure 3(a) presents a comparison of the BLEU scores for the base model, fine-tuned LoRA model, and several configurations in which the key neurons were individually or collectively silenced.
The LoRA fine-tuning process yielded a significant improvement over the base model, increasing the average BLEU score from 0.027 to 0.150. The effects of silencing individual neurons varied. As shown in Figure 3(a), deactivating neuron #890 had no effect on the score (0.150), while silencing neuron #560 led to a marginal improvement (0.152). Most other interventions caused slight performance degradation (e.g., #1066 to 0.147). In contrast, silencing the entire group of six key neurons simultaneously (LoRA-Silenced-Key6) produced the largest performance decrease, reducing the average BLEU score to 0.139.
The LoRA update also induced a significant change in the response style. The base model generated verbose answers with an average length of 71.4 words, whereas the LoRA model produced more concise responses averaging 22.9 words (Figure 3(b)). Silencing interventions did not substantially alter this conciseness.
Figure 3: Comparison of model performance on the Question-Answering task: (a) BLEU Score, (b) Average Answer Length.
4.3 Statistical Significance of Performance Changes
We performed a Wilcoxon signed-rank test to validate these observations. The results are shown in Figure 4. The performance improvement from the Base to the LoRA model was highly statistically significant ($p\ll 0.05$). The performance degradation from silencing the entire key neuron group (LoRA-Silenced-Key6) and the most amplified individual neurons (#1066, #1086) were also statistically significant ($p<0.05$). The changes observed from silencing other individual key neurons did not meet the threshold for statistical significance.
Figure 4: Statistical significance test results (-log10(p-value)). Higher bars indicate greater statistical significance. The red dashed line represents the p=0.05 threshold.
5 DISCUSSION
This section interprets the findings presented in the Results section, focusing on their scientific and practical implications for understanding AI behavior in nuclear safety contexts. First, we discuss the high-level interpretation of our findings, supporting a circuit-based view of knowledge encoding. We then performed a functional analysis of individual key neurons, attributing specific roles to the qualitative results of our silencing experiments. Finally, we directly connect these mechanistic insights to their implications for nuclear safety.
5.1 Interpretation of Key Findings: From Independent Experts to Neural Circuits
Our results provide compelling evidence for a circuit-based hypothesis regarding the encoding of domain expertise in a fine-tuned LLM. The finding that silencing most key neurons individually was not statistically significant while silencing the group collectively was strongly refutes a simple model in which each specialized neuron acts as an independent expert. Instead, it suggests that domain knowledge is a distributed property of a neural circuit, and these neurons act in concert to produce expert-level, domain-specific outputs.
Interestingly, the most amplified neurons (#1066, #1086) were key exceptions, as their individual deactivation was sufficient to cause a statistically significant, albeit small, decrease in performance. This suggests a more nuanced, hybrid model in which expertise is encoded in a distributed circuit, but certain highly influential ”hub neurons” may serve as critical nodes within that circuit. Nevertheless, the fact that the collective silencing had a far greater impact than silencing, even in this single most important neuron, underscores the primary importance of the circuit as a whole.
This nuanced view is further supported by intriguing anomalies observed when silencing other key neurons. For instance, deactivating neuron #890 resulted in no performance change, while deactivating neuron #560 paradoxically led to a slight improvement in the BLEU score. These results suggest a functional differentiation within the specialized neuron group, where individual neurons may play distinct roles such as refinement, suppression of undesirable traits, or context-specific activation, rather than uniformly contributing to knowledge generation.
5.2 Functional Differentiation of Key Neurons
Based on a qualitative analysis of the model’s outputs when specific neurons were silenced, we can hypothesize the distinct roles of these key neurons within the larger circuit.
  • Neurons #1066 & #1086: Key Concept Neurons. As the most influential ”hub neurons”, their function appears to be encoding and retrieval of core, domain-specific technical concepts and entities. When neuron #1066 was silenced, the model’s answer to a question about Swedish SAMG updates lost specific institutional names like ”Swedish Defence Equipment (SDEk)” and defaulted to a generic, uninformative phrase like ”a new version of the SAMG system”. This indicates that these neurons are critical in providing the factual terminological backbone of an expert’s answer.
  • Neuron #941 (Suppressed): The Generalist Language Suppressor. This neuron’s activation was decreased during LoRA fine-tuning, suggesting that it performs a function useful for general language but detrimental to technical accuracy. The Base model often produced conversational, non-committal answers (”You’ve hit on a really important and complex question!…”) and hallucinated incorrect acronyms (”BWR (Bushel-Wall Reactor)”). LoRA learned to suppress neuron #941 to inhibit these undesirable generalist tendencies, allowing more specialized neurons to generate concise, factual responses.
  • Neuron #176: Procedural Logic Neuron. This neuron seems to specialize in encoding the logical relationship between different procedures or states. When asked about the transition from EOP to SAG, silencing #176 caused the model to produce a nonsensical answer about an ””S” sequence”. However, the functional LoRA model correctly identified the condition for the switch. This suggests that neuron #176 is crucial for representing the ”if-then” logic inherent in operational guidelines.
  • Neuron #560: Over-aggressive Generalization Neuron. Paradoxically, silencing this neuron improved the BLEU score. A review of its outputs suggests that it attempts to make answers more concise but sometimes does so by omitting key details. For example, when asked about maintaining water levels for jet pumps, the standard LoRA model stated it was for operating ”effectively and efficiently,” whereas the #560-silenced model more accurately stated it was for operating ”effectively and to prevent damage to the pump.” Silencing this neuron removed a flawed simplification, resulting in a more precise and higher-scoring answer.
  • Neuron #890: Redundant or Context-Specific Neuron. Silencing this neuron had no observable impact on either the BLEU score or qualitative output for any question in the evaluation set. This strongly implies that its role is either fully redundant and compensated for by other neurons in the circuit or that it is highly specialized for a narrow context that was not present in our evaluation data.
5.3 Nuclear Safety Implications of Neuron Silencing
The quantitative performance degradation and functional roles inferred above have direct and significant implications for nuclear safety. The failure modes observed when silencing these key neurons mirror the types of errors that could lead to serious operational incidents. A qualitative analysis of the model outputs revealed three primary failure modes with significant safety consequences, as detailed in Table 2.
Table 2: Comparison of Answer Quality After Silencing Key Neurons
Question Answer from LoRA model (Good) Answer from LoRA-Silenced-Key6 model (Degraded)
What is the primary function of the BWR-UH-HPCI system? The primary function of the BWR-UH-HPCI system is to automatically provide high-pressure makeup water to the reactor vessel to maintain core cooling in the event of a small break Loss of Coolant Accident (LOCA) that does not result in rapid reactor depressurization. The BWR-UH-HPCI system is designed to maintain the water level in the reactor pressure vessel (RPV).
What happens during a station blackout (SBO) event? During an SBO, all offsite and onsite AC power is lost. The High-Pressure Coolant Injection (HPCI) and Reactor Core Isolation Cooling (RCIC) systems, which are steam-driven, become the primary means of maintaining reactor water level. DC power from station batteries is crucial for their control and operation. A station blackout is a loss of power. The reactor will shut down.
Describe the function of the Automatic Depressurization System (ADS). The Automatic Depressurization System (ADS) function is to rapidly reduce reactor pressure during a LOCA, allowing low-pressure core cooling systems like LPCI and LPCS to inject water. It is automatically initiated by signals indicating low reactor water level and high drywell pressure. The ADS reduces pressure in the reactor.
Note: HPCI, High-Pressure Coolant Injection; LOCA, Loss of Coolant Accident; RPV, Reactor Pressure Vessel; SBO, Station Blackout; RCIC, Reactor Core Isolation Cooling; DC, Direct Current; ADS, Automatic Depressurization System; LPCI, Low-Pressure Coolant Injection; LPCS, Low-Pressure Core Spray.
As the table illustrates, the identified neural circuit encodes not only isolated facts, but also the structured, safety-oriented logic that is essential for nuclear operations. The degradation of this circuit translates directly to an increase in operational risk, highlighting the necessity of verifying the integrity of such circuits before deployment.
Furthermore, the phenomenon illustrated in Figure 3, where the average answer length dramatically decreased as the BLEU score substantially increased, demonstrates that the LoRA-adapted model learned to be more succinct and effective, prioritizing factual precision over the verbosity of the base model.
6 Conclusion
This research addressed the critical challenge of ensuring the reliability of LLMs in safety-critical domains like nuclear engineering. We demonstrated a replicable methodology to move beyond treating LLMs as opaque black boxes towards a mechanistic understanding of their function. Specifically, by adapting the Gemma model to the BWR domain using LoRA, we identified a sparse set of specialized neurons. Our causal interventions provided strong evidence for a circuit-based explanation by revealing the complex interplay between differentiated neuron roles and their collective impact. While silencing influential ”hub” neurons like #1066 and #1086 individually caused a statistically significant, albeit small, degradation, this effect was far surpassed by the much larger performance drop from deactivating the entire group collectively. This demonstrates that the model’s full expertise is not merely a sum of independent parts but an emergent property of the neural circuit operating in concert.
The core contribution of this study is a pathway towards achieving nuclear-grade AI assurance. The ability to identify, monitor, and causally test knowledge-bearing neural circuits offers a practical solution to the AI V&V challenge, which has limited the deployment of advanced AI in safety-critical nuclear applications. This approach has profound implications for the nuclear regulatory framework. For instance, by establishing baseline activation maps of key circuits during qualification, as specified in standards like IEEE Std 7-4.3.2 [7], operators and regulators can continuously monitor the AI’s internal reasoning and flag deviations and perform targeted re-validation after plant modifications. This aligns with human factors principles (NUREG-0711 [20]) by making AI recommendations more trustworthy and comprehensible.
However, this study had several limitations. Our reliance on the BLEU score, while justified for measuring output precision, does not verify factual accuracy. Therefore, a key area for future work is the development of domain-specific evaluation metrics that can automatically verify technical correctness against a trusted knowledge base. Furthermore, although we have identified causally important neurons, fully decoding the specific concept that each circuit represents requires more advanced techniques, such as feature visualization.
Future work will extend this methodology to other critical nuclear systems, including Pressurized Water Reactors (PWRs) and advanced reactor designs. To confirm the broader applicability of our findings, it is crucial to validate this approach across different model architectures, model scales, and technical documentation corpora.
The long-term vision is to translate these findings into practical, real-time monitoring protocols suitable for integration with plant digital infrastructure, ensuring that AI serves as a verifiable and trustworthy tool for the future of nuclear energy.
Code Availability
The source code used to generate the results and figures in this paper is publicly available on GitHub at https://github.com/sixticket/llm-xai-nuclear-paper.